OpenStudy (anonymous):

1. Which of the following inequalities matches the graph? graph of an inequality with a solid line through the points (-2, 0) and (0, 3) with shading below the line y < 1 over 3 x + 3 y less than or greater to 3 over 2 x + 3 y greater than or equal to 3 over 2 x + 3 The correct inequality is not listed.

OpenStudy (texaschic101):

a solid line indicates that there is an equal sign in the problem shading below the line indicates a less then sign is used. Now let us find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (-2,0)...x1 = -2 and y1 = 0 (0,3)...x2 = 0 and y2 = 3 now we sub slope(m) = (3 - 0) / (0 - (-2) slope(m) = 3/(0 + 2) slope(m) = 3/2 Now use y = mx + b form...and since we already found the slope(m), we need to find the y intercept(b). y = mx + b slope(m) = 3/2 use either of your points (0,3)...x = 0 and y = 3 now sub and find b 3 = 3/2(0) + b 3 = b so your equation is : y = 3/2x + 3....but remember what I said before, a solid line indicates an equal sign and shading below the line indicates less then. Therefore, your inequality is : y <= 3/2x + 3

OpenStudy (anonymous):

Its choice B?

OpenStudy (anonymous):

@texaschic101